Q1: If the recorded monthly wind speeds recorded in Bahrain as follows: 4.90, 3.8, 6.9, 7.0, 2.9, 4.5, 5.0, 5.3, 5.8, 6.5, 6.8, 4.9, ms1, then estimate the following: 1. The potential of the wind power The maximum extractable electrical power if the efficiency of the wind generator is 35% and the blade size of the wind generator is 5 m and the air density is 1.25 kg m-3 2.
Answer #1

ANSWER-

1. POTENTIAL OF THE WIND POWER =/2fCV3-1/2 x 1.23 x .4 × (5.36) = 37.88

                                                                      =37.88 W/M2

2. POWER=1/2f0V3A 1/2x 1.23 × .4 × (5.36)3 × 22/7 × 25 = 2976 .5W

SO, MAXIMUM EXTRACTABLE POWER =35% OF 2976.5W

                                                                      = 1041.79W